OpenStudy (anonymous):

How can you show that S={v_1,v_2} is a linearly independent set in R^3 when you're only given 2 vectors?

OpenStudy (anonymous):

(1, -1,0) and (3,1,-2) I rref them and got

OpenStudy (anonymous):

>> [1 0 -.5; 0 1 -.5]

OpenStudy (anonymous):

guess you have to show that if \[av_1+bv_2=0\] then \[a=b=0\] yes?

OpenStudy (anonymous):

Yeah so would I get x_1= -5 and x_2 = -.5?

OpenStudy (anonymous):

-x_1+x_2 = 0? so hence its linear independent?

OpenStudy (anonymous):

i get the equations \[a+3b=0\] \[-a+b=0\] \[-2b=0\]

OpenStudy (anonymous):

last one says b=0 and since -a+b=0 this means a = 0 as well

OpenStudy (anonymous):

really in this case it just means that one is not a multiple of the other, which it isn't

OpenStudy (anonymous):

Did you row reduce to get those equations?

OpenStudy (anonymous):

no i just did it the donkey way

OpenStudy (anonymous):

\[a(1,-1,0)+b(3,1,-2)=(0,0,0)\]

OpenStudy (anonymous):

I'm not quite following.. but it could just be me..

OpenStudy (anonymous):

I row reduced and got [1 0 -.5; 0 1 -.5]?